# Leveraging temporal features of the divergence measure to detect chaos in conservative systems

### v1

@@ Appendix B Robustness with respect to choices
Appendix C Applications to two stroboscopic maps computed from resonant Hamiltonians
@@
C.1. Results for a modulated pendulum
@@
corresponding to a pendulum model with variable length. For this model, $T=2\pi/\epsilon$ which is large when $\epsilon\ll 1$. We selected $\alpha=0.25$ and $\epsilon=0.1$. The phase space of $P$ together with the temporal laws of $DIV$ on regular and chaotic components are shown in Fig. 12. We found $\gamma_{\textrm{reg}.}=-1.06$ and $\gamma_{\textrm{cht}.}=-0.40$
@@
Table 3. Exponents for the power-laws of $\langle DIV\rangle$ found for the Poincaré map associated to the continuous model $\mathcal{J}$ of Eq. (17) under different dynamical settings. The denomination “observable II” and “observable $\phi$” is an abuse of notation and refer to the discrete variables of the Poincaré mapping associated to the original momentum II and angle $\phi$. The parameters according to time-delay phase space reconstruction Takens, (1981) are $d=5$ and $\tau=1.$
Figure 12. (Left) Phase portrait of the modulated pendulum. The light blue box depicts the initial conditions of 200 members of the ensemble used to estimate the power-law of DIVDIV vs NN. The uniform distribution in the highlighted domain provides roughly the same number of regular and chaotic orbits. (Right) Time dependence of measure DIVDIV for chaotic (red) and regular (blue) parts of the ensemble. Initial conditions of trajectories are shown in left panel.
C.2. Results for two resonances that overlap
(18) $\mathcal{K}(I,\phi,t)=\frac{I^{2}}{2}-(\alpha_{1}\cos(\phi-t)+\alpha_{2}\cos(\phi+t)),$
containing two cat-eye resonances centered at I=1I=-1 and $I=1$. For this model, $T=2\pi$. We choose $\alpha_{1}=0.2$ and $\alpha_{2}=0.2$ (both resonances have equal dynamical weights). The phase space of $P$ together with the temporal laws of $DIV$ on regular and chaotic components are shown in Fig. 13. We founds $\gamma_{\textrm{reg}.}=-1.02$ and $\gamma_{\textrm{cht}.}=-0.56$
$\gamma_{\textrm{reg.}}$ $\gamma_{\textrm{cht.}}$
Original 22D phase space -1.02 -0.52
Observable $I$, reconstruction -1.06 -0.62
Observable $\phi$, reconstruction -1.03 -0.62
Observable $I$, no reconstruction -1.16 -0.59
Observable $\phi$, no reconstruction -1.06 -0.59
Table 4. Same as in Tab. 3 for the continuous model $\mathcal{K}$ of Eq. (18)
Figure 13. (Left) Phase portrait for the resonance overlap Hamiltonian (18). (Right) Similar to Fig. 12.

### v2

@@ Appendix A $DIV$ versus $N$ for ARMA process
Appendix B Applications to two stroboscopic maps computed from resonant Hamiltonians
@@
B.1. Results for two resonances that overlap
We consider the model
(19) $\mathcal{K}(I,\phi,t)=\frac{I^{2}}{2}-(\alpha_{1}\cos(\phi-t)+\alpha_{2}\cos(\phi+t))$
discussed in aMo22 (chapter $6$). For this model, $T=2\pi$. The 2-DoF autonomous counterpart of Eq. (19) reads
(20) $\mathcal{K}=\frac{I^{2}}{2}+J-(\alpha_{1}\cos(\phi-\tau)+\alpha_{2}\cos(\phi+\tau)).$
When $\alpha_{1}=0$ or $\alpha_{2}=0$, we recover the unperturbed integrable Hamiltonian of the pendulum using an ad-hoc canonical change of variables. The phase space contains the single cat-eye resonance centered around either $c_{1}(I)=-1$ or $c_{2}(I)=1$, with half-widths of the separatrices $\delta_{2}=2\sqrt{\alpha_{2}}$ or $\delta_{1}=2\sqrt{\alpha_{1}}$ respectively. Whenever both $\alpha_{1}$ and $\alpha_{2}$ are different from zero, the resonances might overlap and chaos appears. The resonance overlap parameter, also called stochasticity parameter (bCh79), reads
(21) $s=\frac{\delta_{1}+\delta_{2}}{|c_{2}-c_{1}|}=\sqrt{\alpha_{1}}+\sqrt{\alpha_{2}}.$
In our numerical setting, we assign to each resonant eye the same dynamical weight with $\alpha_{1}=\alpha_{2}=1/5$, leading to $s\approx 0.89$. As $s$ is close to $1$, the resonances overlap significantly and macroscopic chaos is expected.
The phase space obtained by iterating $P$ together with the temporal laws of $DIV$ on regular and chaotic components are shown in Fig. 14. The fits of the spatial averages lead to $\gamma_{\textrm{reg}.}=-1.02$ and $\gamma_{\textrm{cht}.}=-0.56$ in the original two-dimensional phase space of $P$.
Figure 14. (Left) Phase portrait for the resonance overlap Hamiltonian (19) obtained with its associated Poincaré map. (Right) The spatial averages <DIV><DIV> follow distinct power laws on the regular and chaotic components.
Although we do not show the various plots here, Tab.  3 summarises the exponents we found under various dynamical setting as we did in Sec. 3, i.e.,by computing the $DIVs$ from observables, including or not the phase-space reconstruction. The exponents found are in accordance with the exponents determined so far.
$\gamma_{\textrm{reg.}}$ $\gamma_{\textrm{cht.}}$
Original 22D phase space -1.02 -0.52
Observable $I$, reconstruction -1.06 -0.62
Observable $\phi$, reconstruction -1.03 -0.62
Observable $I$, no reconstruction -1.16 -0.59
Observable $\phi$, no reconstruction -1.06 -0.59
Table 3. Exponents for the power laws of $\langle DIV\rangle$ found for the Poincaré map associated to the continuous model $\mathcal{K}$ of Eq. (19) under different dynamical settings. The denomination “observable II” and “observable $\phi$” is an abuse of notation to refer to the discrete variables of the Poincaré mapping associated to the original momentum II and angle $\phi$. The parameters according to time-delay phase space reconstruction (fTa81) are $d=5$ and $\tau=1.$
B.2. Results for a modulated pendulum
@@
corresponding to a pendulum model with variable length. This Hamiltonian is discussed in aMo22 (chapter $9$). For this model, $T=2\pi/\epsilon$ which is large when $\epsilon\ll 1$. As the time variable $\tau=\epsilon t$ is slow ($\dot{\tau}=\epsilon$, $\epsilon\ll 1$), this model is paradigmatic of slow chaos where $3$ resonances are $\epsilon$ apart. In fact, using trigonometrical identities, $\mathcal{J}$ might be rewritten as
(23) $\mathcal{J}=\frac{I^{2}}{2}+\epsilon J-\big(\cos\phi+\frac{\alpha}{2}\cos(\phi-\tau)+\frac{\alpha}{2}\cos(\phi+\tau)\big).$
The $3$ harmonics are now clearly apparent. By Hamilton’s canonical equations, one sees that the centres of the resonances $\dot{\phi}=0$, $\dot{\phi}-\dot{\tau}$ and $\dot{\phi}+\dot{\tau}$ correspond respectively to the actions values $I=0$, $I=\epsilon$, $I=-\epsilon$. We selected $\alpha=0.25$ and $\epsilon=0.1$. The phase space of $P$ together with the temporal laws of $DIV$ on regular and chaotic components are shown in Fig. 15. We found $\gamma_{\textrm{reg}.}=-1.06$ and $\gamma_{\textrm{cht}.}=-0.40$.
Figure 15. (Left) Phase portrait of the modulated pendulum. The light blue box depicts the initial conditions of 200 members of the ensemble used to estimate the power law of DIVDIV vs NN. The uniform distribution in the highlighted domain provides roughly the same number of regular and chaotic orbits. (Right) Time dependence of measure DIVDIV for chaotic (red) and regular (blue) parts of the ensemble. Initial conditions of trajectories are shown in left panel.
The exponents found when starting from the knowledge of observables, including the reconstruction of the phase space or not, are presented in Tab. 4.
@@
Table 4. Same as Tab. 3 for the Hamiltonian model of Eq. 22.